7 years ago I was 11 times as old as my daughter. Now I am 4 times as old as her. How old am I?
Answer by checkley75(3666)

x-7=11d
x=4(d+7)
4(d+7)-7=11d
4d+28-7=11d
11d-4d=21
7d=21
d=3 age of daughter 7 years ago. now she is 3+7=10 years old
dad is
x-7=11*3
x-7+33
x=33+7
x=40 years old for the dad now
proof
40=4(3+7)
40=4*10
40=40